# Can GPT Redefine Medical Understanding? Evaluating GPT on Biomedical Machine Reading Comprehension

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable performance on many tasks in different domains. However, their performance in closed-book biomedical machine reading comprehension (MRC) has not been evaluated in depth. In this work, we evaluate GPT on four closed-book biomedical MRC benchmarks. We experiment with different conventional prompting techniques as well as introduce our own novel prompting method. To solve some of the retrieval problems inherent to LLMs, we propose a prompting strategy named Implicit Retrieval Augmented Generation (RAG) that alleviates the need for using vector databases to retrieve important chunks in traditional RAG setups. Moreover, we report qualitative assessments on the natural language generation outputs from our approach. The results show that our new prompting technique is able to get the best performance in two out of four datasets and ranks second in rest of them. Experiments show that modern-day LLMs like GPT even in a zero-shot setting can outperform supervised models, leading to new state-of-the-art (SoTA) results on two of the benchmarks.
@@ 1 Introduction
Machine Reading Comprehension (MRC) is defined as a task where a system tries to answer a question based on a given context. The context could be anything ranging from a couple of passages to a list of documents. Even though there has been a lot of research done on MRC, there are still several challenges that are faced when dealing with MRC tasks Sugawara et al. (2022) such as inability to take handle long-range dependencies when reasoning and domain adaption. Recent improvements in large language modeling has alleviated a lot of the aforementioned issues.
MRC in the biomedical domain Hermann et al. (2015); Baradaran et al. (2022) has remained a long-standing area of research. Sun et al. (2022) proposed a novel tagging strategy for the joint extraction of biomedical entities and relations in the MRC framework. Solving a biomedical MRC task faces various challenges including large intricate in-domain vocabulary, dependency on global knowledge, etc. Due to these challenges, there is a wide gap between the performance of conventional methods in the general domain and that of the biomedical domain. Although, traditional machine learning models did show some improvement, they have never been any close to human baselines or gold standards. Contrary to this preconceived notion, modern-day LLMs have shown remarkable performance on many biomedical tasks Nori et al. (2023); Yang et al. (2023); Cheng et al. (2023).
An open-book setting requires the LLMs to not only rely on the given context but also use the global knowledge embedded in their parameters to answer a query. On the other hand, in a closed-book setting, the LLMs rely solely on the given context to answer a query. Some of the datasets corresponding to open-book are Zhang et al. (2018); Pal et al. (2022). Similarly, Berant et al. (2014b); Pappas et al. (2018); Zhu et al. (2020) comprise of the datasets for closed-book. Recent LLMs have attained unprecedented performance in a wide array of natural language processing (NLP) tasks Chang et al. (2023). Although their performance has been evaluated on a multitude of MRC benchmarks in an open-book setting, their performance in a closed-book setting has been understudied. In this work, we fill out this missing gap by evaluating GPT OpenAI (2023) on standard closed-book MRC benchmarks of the biomedical domain. The key contributions are:
1. We evaluate different prompting techniques with GPT on four closed-book biomedical MRC benchmarks and report new SoTA results.
@@ 2 Related Work
In order to elicit an answer from an LLM like GPT, one needs to prompt it in natural language in an optimal manner to retrieve the intended answer. To that end, there has been tremendous development in finding optimal methods for prompting LLMs. The maximum performance boost has been seen from the Chain-of-Thought (CoT) Reasoning Wei et al. (2022) prompting strategy which asks the LLM to explain how it arrived at the answer. More recently, Analogical Reasoning (AR) Yasunaga et al. (2023a) has been proposed that achieve drastically better performance than CoT and other prompting techniques. AR works by asking the LLM to reason about a problem by giving analogies which in return forces the model to leverage the global knowledge encoded in it. While prompting methods like CoT and AR improve LLM’s performance by exploiting the model’s global knowledge embedded in it, there has been an increase in developing novel techniques, especially for cases where the context that needs to be searched through to answer the asked query is huge. The context could be one huge document or a combination of multiple short/long documents. In such scenarios, it is very important to identify only the relevant chunks of the context required for the underlying task and pay attention to them. These emerging methods come under the umbrella of Retrieval Augmented Generation (RAG) Lewis et al. (2020), which has been shown to improve the performance of LLMs by retrieving contextually relevant information from corpora. The basic methodology behind RAG is to use, embed, and store context in a vector database. These embeddings can then be retrieved based on their semantic similarity to the query.
All the aforementioned prompting methods help interface and contextualize inputs in a better manner for LLMs. While the efficacy of these methods has been seen on several large benchmarks in different domains, however, the degree to which they help in closed-book biomedical MRC has been understudied. Mahbub et al. (2022) presented an adversarial learning-based domain adaptation framework for the biomedical MRC task to address the discrepancies in the marginal distributions between the datasets of the general and biomedical domains. Nori et al. (2023) evaluated GPT on medical competency examinations and benchmark datasets. Even though their work talks about biomedical MRC, it concentrates only on open-book benchmarks. Similarly, Singhal et al. (2023) evaluated Med-PaLM 2 on medical competency examinations and thus focuses on open-book biomedical MRC.
@@ 3 Datasets
The four datasets from the biomedical and healthcare domain we choose to explore and analyze the performance of GPT are ProcessBank Berant et al. (2014b), BioMRC Pappas et al. (2020), MASH-QA Zhu et al. (2020) and CliCR Šuster and Daelemans (2018). There are a multiple reasons for selecting these four datasets. First, we want to focus on datasets that have not yet been evaluated by modern-day LLMs like GPT. Next, we want to pick up datasets that vary in their statistics and nature. Finally, based on our understanding, these 6 datasets covered the majority of research around closed-book MRC in the biomedical field.
@@ 5 Results & Analysis
4. There are a lot of entities which are semantically and syntactically same but still belong to different ontologies and thus have different entity IDs. For example, in one case, GPT generates the answer amino acids where the correct answer is amino acid but since both of these entities have different IDs, this answer had to be marked wrong.
@@
Another important aspect to note here apart from the lack of structure in the dataset is the overall system design of supervised models which are being compared to GPT when talking about SoTA. Supervised models use 70%-80% of the available data as their training set which allows their parameters to get a good idea about the nuisances of the dataset whereas in case of GPT, all our experiments are being conducted in a zero-shot setting. Also since GPT is a generative model, the chances of GPT generating an answer not present in the candidate answer list despite the final answer being semantically and syntactically the same is really high. But a supervised model is never going to face this problem as it makes its prediction based on the confidence score for each candidate answer and thus the final answer is always going to be present in the candidate answer list.
@@
Bert 0.9 0.25 0.56 0.16
RoBERTa 0.9 0.29 0.58 0.19
XLNet 0.9 0.29 0.56 0.20
@@
2. There is a correlation observed between increase in the number of sections and depreciating performance of Implicit RAG. The reason is that the answers in this dataset are long span and thus with increasing number of sections, there is a loss of contextual continuity as the ground truth answers can get split across multiple sections. This ends up confusing the LLM making it difficult to choose the right set of sentences from different sections. Hence, it performs the best when asked to extract just one section from the context. But because MASH-QA contains answers which can be present in disjoint spans of the context, extracting just one section is not able to make Implicit RAG the best performing prompting method.
@@
3. One question which may arise is whether extracting one section or having the hyper-parameter number of sections set to 1 for Implicit RAG makes it same as that of Basic prompting. Implicit RAG and Basic prompting become same only when we are not only extracting just one section from the context but also when the hyper-parameter number of words for Implicit RAG is set equal to the length of the entire given context. But for MASH-QA, the hyper-parameter number of words is set to 300 with the lower limit being 0 and upper limit being 300 and hence they are different.
@@
1. The authors of CliCR mention that for the training of supervised models, only those instances are used for which at least one ground-truth answer from the set of ground-truth answers occurs in the clinical case report or the context. But for the evaluation part, for both validation and test sets even those datapoints are included where there is no intersection between ground-truth set and the entities mentioned in the context. This favors supervised learning settings as supervised models have a separate training and development set which can allow the parameters of the model to learn such cues. GPT is still able perform better possibly because the global knowledge embedded in its parameters gives it enough evidence to perform well.
2. The authors of CliCR compare various skills in their work between the then SoTA (GPT is the new SoTA) model GA-NoEnt and Human Expert and show that there still exists a huge gap between them. Since GPT is able to achieve almost Human Expert level performance, we can expect GPT to show similar capabilities in other MRC tasks.

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable performance on many tasks in different domains. However, their performance in contextual biomedical machine reading comprehension (MRC) has not been evaluated in depth. In this work, we evaluate GPT on four contextual biomedical MRC benchmarks. We experiment with different conventional prompting techniques as well as introduce our own novel prompting method. To solve some of the retrieval problems inherent to LLMs, we propose a prompting strategy named Implicit Retrieval Augmented Generation (RAG) that alleviates the need for using vector databases to retrieve important chunks in traditional RAG setups. Moreover, we report qualitative assessments on the natural language generation outputs from our approach. The results show that our new prompting technique is able to get the best performance in two out of four datasets and ranks second in rest of them. Experiments show that modern-day LLMs like GPT even in a zero-shot setting can outperform supervised models, leading to new state-of-the-art (SoTA) results on two of the benchmarks.
@@ 1 Introduction
Machine Reading Comprehension (MRC) is defined as a task where a system tries to answer a question based on a given context. The context could be anything ranging from a couple of passages to a list of documents. Even though much research has been conducted on MRC, several challenges remain when dealing with MRC tasks Sugawara et al. (2022), such as the inability to handle long-range dependencies when trying to do reasoning and domain adaptation. Recent improvements in large language modeling has alleviated a lot of the aforementioned issues.
MRC in the biomedical domain Hermann et al. (2015); Baradaran et al. (2022) has always been a key area of research. Solving a biomedical MRC task faces various challenges including large intricate in-domain vocabulary, dependency on global knowledge, etc. Due to these challenges, there is a wide gap between the performance of conventional methods in the general domain and that of the biomedical domain. Although, traditional machine learning models did show some improvement, they have never been any close to human baselines or gold standards. Contrary to this preconceived notion, modern-day LLMs have shown remarkable performance on many biomedical tasks Nori et al. (2023); Yang et al. (2023); Cheng et al. (2023).
MRC can have different variations in itself. A contextual MRC requires the LLMs to answer a query solely by relying on a given context. In contrast, a context-free MRC relies on model’s embedded knowledge base or any open-source knowledge base, such as Wikipedia, to answer a query instead of using only the context provided. Some of the datasets corresponding to context-free MRC are Zhang et al. (2018); Pal et al. (2022). These datasets are classified under context-free MRC because the given context is not sufficient to answer the questions. There is a definite need to explore the LLM’s inherent knowledge base or any other source of knowledge to answer these queries. Similarly, Berant et al. (2014b); Pappas et al. (2018); Zhu et al. (2020) comprise of the datasets for contextual MRC. Again, these datasets have been categorized under contextual MRC because the queries asked can be correctly answered just by looking at the provided context. There is no need to induce any kind of external knowledge in order to answer these questions.
Recent LLMs have attained unprecedented performance in a wide array of natural language processing (NLP) tasks Chang et al. (2023); Vatsal and Dubey (2024). Although their performance have been evaluated on a multitude of MRC benchmarks in a context-free setting, their performance in a contextual setting has been understudied. In this work, we fill out this missing gap by evaluating GPT OpenAI (2023) on standard contextual MRC benchmarks of the biomedical domain. The key contributions are:
1. We evaluate different prompting techniques with GPT on four contextual biomedical MRC benchmarks and report new SoTA results.
@@ 2 Related Work
In order to elicit an answer from an LLM like GPT, one needs to prompt it in natural language in an optimal manner to retrieve the intended answer. To that end, there has been tremendous development in finding optimal methods for prompting LLMs. The maximum performance boost has been seen from the Chain-of-Thought (CoT) Reasoning Wei et al. (2022) prompting strategy which asks the LLM to explain how it arrived at the answer. More recently, Analogical Reasoning (AR) Yasunaga et al. (2023a) has been proposed that achieves drastically better performance than CoT and other prompting techniques. AR works by asking the LLM to reason about a problem by giving analogies which in return forces the model to leverage the global knowledge encoded in it. While prompting methods like CoT and AR improve LLM’s performance by exploiting the model’s global knowledge embedded in it, there has been an increase in developing novel techniques, especially for cases where the context that needs to be searched through to answer the asked query is huge. The context could be one huge document or a combination of multiple short/long documents. In such scenarios, it is very important to identify only the relevant chunks of the context required for the underlying task and pay attention to them. These emerging methods come under the umbrella of Retrieval Augmented Generation (RAG) Lewis et al. (2020), which has been shown to improve the performance of LLMs by retrieving contextually relevant information from corpora. The basic methodology behind RAG is to use, embed, and store context in a vector database. These embeddings can then be retrieved based on their semantic similarity to the query.
All the aforementioned prompting methods help interface and contextualize inputs in a better manner for LLMs. While the efficacy of these methods has been seen on several large benchmarks in different domains, however, the degree to which they help in contextual biomedical MRC has been understudied. Mahbub et al. (2022) presented an adversarial learning-based domain adaptation framework for the biomedical MRC task to address the discrepancies in the marginal distributions between the datasets of the general and biomedical domains. Nori et al. (2023) evaluated GPT on medical competency examinations and benchmark datasets. Even though their work talks about biomedical MRC, it concentrates only on context-free MRC benchmarks. Similarly, Singhal et al. (2023) evaluated Med-PaLM 2 on medical competency examinations and thus focuses on context-free biomedical MRC.
@@ 3 Datasets
The four datasets from the biomedical and healthcare domain we choose to explore and analyze the performance of GPT are ProcessBank Berant et al. (2014b), BioMRC Pappas et al. (2020), MASH-QA Zhu et al. (2020) and CliCR Šuster and Daelemans (2018). There are a multiple reasons for selecting these four datasets. First, we want to focus on datasets that have not yet been evaluated by modern-day LLMs like GPT. Next, we want to pick up datasets that vary in their statistics and nature. Finally, based on our understanding, these 4 datasets covered the majority of research around contextual MRC in the biomedical field.
@@ 5 Results & Analysis
4. There are a lot of entities which are semantically and syntactically the same but still belong to different ontologies and thus have different entity IDs. For example, in one case, GPT generates the answer amino acids where the correct answer is amino acid but since both of these entities have different IDs, this answer had to be marked wrong.
@@
Another important aspect to note here apart from the lack of structure in the dataset is the overall system design of supervised models which are being compared to GPT when talking about SoTA. Supervised models use 70%-80% of the available data as their training set which allows their parameters to get a good idea about the nuances of the dataset whereas in case of GPT, all our experiments are being conducted in a zero-shot setting. Also since GPT is a generative model, the chances of GPT generating an answer not present in the candidate answer list despite the final answer being semantically and syntactically the same is really high. But a supervised model is never going to face this problem as it makes its prediction based on the confidence score for each candidate answer and thus the final answer is always going to be present in the candidate answer list.
@@
Bert 0.09 0.25 0.56 0.16
RoBERTa 0.09 0.29 0.58 0.19
XLNet 0.09 0.29 0.56 0.20
@@
2. There is a correlation observed between increase in the number of sections and decreasing performance of Implicit RAG. The reason is that the answers in this dataset are long span and thus with increasing number of sections, there is a loss of contextual continuity as the ground truth answers can get split across multiple sections. This ends up confusing the LLM making it difficult to choose the right set of sentences from different sections. Hence, it performs the best when asked to extract just one section from the context. But because MASH-QA contains answers which can be present in disjoint spans of the context, extracting just one section is not able to make Implicit RAG the best performing prompting method.
@@
3. One question which may arise is whether extracting one section or having the hyper-parameter number of sections set to 1 for Implicit RAG makes it same as that of Basic prompting. Implicit RAG and Basic prompting become the same only when we are not only extracting just one section from the context but also when the hyper-parameter number of words for Implicit RAG is set equal to the length of the entire given context. But for MASH-QA, the hyper-parameter number of words is set to 300 with the lower limit being 0 and upper limit being 300 and hence they are different.
@@
1. The authors of CliCR mention that for the training of supervised models, only those instances are used for which at least one ground-truth answer from the set of ground-truth answers occurs in the clinical case report or the context. But for the evaluation part, for both validation and test sets even those datapoints are included where there is no intersection between ground-truth set and the entities mentioned in the context. This favors supervised learning settings as supervised models have a separate training and development set which can allow the parameters of the model to learn such cues. GPT is still able to perform better possibly because the global knowledge embedded in its parameters gives it enough evidence to perform well.
2. The authors of CliCR compare various skills in their work between the previous SoTA (GPT is the new SoTA) model GA-NoEnt and Human Expert and show that there still exists a huge gap between them. Since GPT is able to achieve almost Human Expert level performance, we can expect it to show similar capabilities in other MRC tasks.